Sol Summary – December 27th

Crew 236 Sol Summary Report 27-DEC-2021

Sol: 8

Summary Title: Three Cheers for Mancos Shale!

Author’s name: Kasey Hilton

Mission Status: Nominal

Sol Activity Summary: The day started with an early wake up so the crew could get out the door for an EVA. Executive officer, Cesare, has been chasing a kind of rock called a Mancos shale for the last four years and we were going to get it today on EVA if it was the last thing we did. And we did just that! Cesare’s face lit up like a Martian Christmas tree as we entered the area near Copernicus Hwy that held the precious rock. After returning to the Hab, just about everyone knocked out for an afternoon nap. After waking up, I continued working on my outreach projects, which includes answering questions for an elementary school class and making a MDRS cookbook. We’re ending the night with fresh-out-of-the-oven brownies and Apollo 13.

Look Ahead Plan: EVA scheduled for tomorrow afternoon/morning. There’s a small chance of snow in the afternoon!

Anomalies in work: No anomalies in work

Weather: Partly cloudy, windy, temperatures in the 30s

Crew Physical Status: Nominal

EVA: EVA to the moon overlook and the Morrison and Mancos shale boundary was a success. Multiple samples of gypsum and Mancos shale were collected for a research project.

Reports to be filed: Operations Report, GreenHab Report, EVA Report, EVA Request, Health and Safety Report, Sol Summary, and Journalist Report (with photos)

Support Requested: None

Operations Report – December 27th

Crew 236 Operations Report 27-DEC-2021

SOL: 8

Name of person filing report: Pavithra Ravi

Non-nominal systems: None


Spirit rover used: Yes

Hours (before EVA): 146.9

Beginning charge (Before EVA): 100%

Ending charge (On return from EVA, before recharging): 64%

Currently charging: Yes

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: Yes

Hours: 156.3

Beginning charge: 100%

Ending charge: 71%

Currently charging: Yes

Perseverance rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

General notes and comments: The front door got stuck in the wood and
had to be freed with a crowbar. The outer door has once again been
secured with a rope (held in place by the closed inner door).

The toilet has been successfully repaired (thank you!).

Summary of Hab operations: Nominal


Water (static tank): approx. 239 gallons

Water (loft tank): 54 gallons

Water Meter: 01541216 units

Static to Loft Pump used – yes or no: Yes. It was filled from 28 to 54
gallons at about 19:30 today.

Static tank pipe heater (on or off): On

Static tank heater (on or off) Off

Toilet tank emptied (no or yes): No

Summary of internet: Nominal

Summary of suits and radios: Nominal

Summary of GreenHab operations: Nominal

WATER USE: 10 gallons were used today. This leaves about 272 gallons
in the tank.

Heater (on or off): On

Supplemental light (on or off): On

Harvest: Cucumber (132 g) and Basil (less than 1 g)

Summary of ScienceDome operations: Nominal

Dual split, Heat or AC (on or off): off

Summary of RAM operations: Tools used, work done

Nothing used. Everything taken (zipties, wrench, Teflon tape) from the
RAM has been returned.

Summary of any observatory issues: Nothing new to report.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support:


EVA Report – December 27th

Crew 236 EVA Report 27-DEC-2021

EVA # 8

Author: Kasey Hilton

Purpose of EVA: Collection of samples from region D and Mancos Shales from Copernicus Hwy for analysis of water content and analog for rocks found in Gale Crater

Start time: 9:00

End time: 13:00

Narrative: EVA prep started around 8:30 and the group going on EVA was in the airlock at 9:02. Shortly after, the two pairs climbed in Spirit and Curiosity and headed towards region D using Cow Dung Rd and Brahe Hwy. About 30 minutes later, the group reached the moon overlook (region D) and parked the rovers to start walking down Copernicus Hwy in search of gypsum and Mancos shale. At the intersection of Brahe Hwy and Copernicus Hwy, the group found multiple samples of gypsum. Further down Copernicus Hwy and closer to the boundary walls, the group found samples of Mancos shales. Cesare (XO) was particularly excited about the mancos shale because he had been trying to find some for the last four years without any luck. The group turned around and walked back along Copernicus Hwy until they reached the rovers. They arrived back at the Hab at 11:46, happy that the research objectives they set out to accomplish had been achieved.

Destination: Region D and the Morrison and Mancos Shale boundary on Copernicus (2 stops)

Coordinates (use UTM NAD27 CONUS): N4254000, W515575

Participants: Kasey (CMD), Cesare (XO), Dylan (ASTRO), Pavi (ENG)

Road(s) and routes per MDRS Map: Cow Dung Rd north to Brahe Hwy, Brahe Hwy west to Copernicus Hwy, south on Copernicus Hwy

Mode of travel: Driving to the west side of Region D and walking down Copernicus Hwy

Greenhab Report – December 27th

GreenHab Officer: Vladimir Zeltsman

Environmental Control: Around 17% Humidity

Average Temperatures: 78degrees

Hours of Supplemental Light: 7 hours

Daily Water Usage For Crops: 10 gallons

Daily water Usage For Research and/or Other Purposes: None

Water In Blue Tank: 172 Gallons

Times Of Watering For Crops: 10 am

Changes to Crops: Cut dead leaves of the cucumbers and zucchinis and sprayed the leaves with soapy water

Narrative: small spider mite infestation that I am taking care of. It seems that they are only on the cucumbers and zucchinis.

Harvest: 132g Cucumber, <1 g basil

Support/supplies needed: None

Copyright © The Mars Society. All rights reserved. | Main Site